# Consistency of variational inference for Besov priors in non-linear inverse problems

### v1

@@ 5.3. Contraction rates for high-dimensional Besov priors
where $\theta_{lr}=0$ for all $l>J$. Using (1.3) and wavelet characterization of Sobolev norms (see [15, section 4]), we deduce that
@@ 6.1. Some facts about the univariate p-exponential distribution
We first calculate the normalizing constant $C_{p,b}$:
@@ 6.2. Concentration of Besov priors
Consider $\Pi^{\prime}$ defined in Definition 2.2. Then,for $p\in[1,2]$, any $h\in B_{pp}^{\alpha}(\mathbb{R}^{d})$ and any symmetric convex Borel-measurable $A\subset L_{2}(\mathbb{R}^{d})$,it holds
@@
Consider $\Pi^{\prime}_{J}$ defined in Definition 2.3. Then, for $p\geq 1$, any $h\in B_{pp}^{\alpha}(\mathbb{R}^{d})$ and any symmetric convex Borel-measurable $A\subset L_{2}(\mathbb{R}^{d})$, there exists a constant $c>0$ such that
$\Pi^{\prime}_{J}(h+A)\geq\big{(}\frac{1}{2}\big{)}^{c2^{Jd}}e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{B_{pp}^{\alpha}}}\Pi^{\prime}_{J}(A).$
@@
$\Pi^{\prime}_{J}(A+h)=\int_{A}\exp\Big{\{}{\sum_{l=-1}^{J}\sum_{r\in R_{l}}\Big{(}V\big{(}\frac{\theta_{lr}}{\gamma_{lr}}\big{)}-V\big{(}\frac{\theta_{lr}-h_{lr}}{\gamma_{lr}}\big{)}\Big{)}}\Big{\}}\Pi^{\prime}_{J}(d\theta)$
$=$ $e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{\tilde{\mathcal{Z}}}}\int_{A}\exp\Big{\{}{\sum_{l=-1}^{J}\sum_{r\in R_{l}}\Big{(}V\big{(}\frac{\theta_{lr}}{\gamma_{lr}}\big{)}+V\big{(}\frac{h_{lr}}{\gamma_{lr}}\big{)}-V\big{(}\frac{\theta_{lr}-h_{lr}}{\gamma_{lr}}\big{)}\Big{)}}\Big{\}}\Pi^{\prime}_{J}(d\theta)$
$=$ $e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{\tilde{\mathcal{Z}}}}\int_{A}\exp\Big{\{}\sum_{l=-1}^{J-1}\sum_{r\in R_{l}}\Big{(}V\big{(}\frac{\theta_{lr}}{\gamma_{lr}}\big{)}+V\big{(}\frac{h_{lr}}{\gamma_{lr}}\big{)}-V\big{(}\frac{\theta_{lr}-h_{lr}}{\gamma_{lr}}\big{)}\Big{)}$
$\qquad\qquad\qquad+\sum_{r\in R_{l}/\{\tilde{r}\}}\Big{(}V\big{(}\frac{\theta_{Jr}}{\gamma_{Jr}}\big{)}+V\big{(}\frac{h_{Jr}}{\gamma_{Jr}}\big{)}-V\big{(}\frac{\theta_{Jr}-h_{Jr}}{\gamma_{Jr}}\big{)}\Big{)}\Big{\}}$
@@
$\Pi^{\prime}_{J}(A+h)\geq e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{\tilde{\mathcal{Z}}}}\int_{A}\frac{1}{2}\exp\Big{\{}\sum_{l=-1}^{J-1}\sum_{r\in R_{l}}\Big{(}V\big{(}\frac{\theta_{lr}}{\gamma_{lr}}\big{)}+V\big{(}\frac{h_{lr}}{\gamma_{lr}}\big{)}-V\big{(}\frac{\theta_{lr}-h_{lr}}{\gamma_{lr}}\big{)}\Big{)}$
@@
$\geq$ $e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{\tilde{\mathcal{Z}}}}\int_{A}\big{(}\frac{1}{2}\big{)}^{\lvert{R_{J}}\rvert}\exp\Big{\{}{\sum_{l=-1}^{J-1}\sum_{r\in R_{l}}\Big{(}V\big{(}\frac{\theta_{lr}}{\gamma_{lr}}\big{)}+V\big{(}\frac{h_{lr}}{\gamma_{lr}}\big{)}-V\big{(}\frac{\theta_{lr}-h_{lr}}{\gamma_{lr}}\big{)}\Big{)}}\Big{\}}\Pi^{\prime}_{J}(d\theta)$
$\geq$ $\big{(}\frac{1}{2}\big{)}^{\sum_{l=-1}^{J}\lvert{R_{l}}\rvert}e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{\tilde{\mathcal{Z}}}}\Pi^{\prime}_{J}(A)\geq\big{(}\frac{1}{2}\big{)}^{c2^{Jd}}e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{\tilde{\mathcal{Z}}}}\Pi^{\prime}_{J}(A).$
@@
We note that Lemmas 6.2, 6.5, 6.4, and 6.7 can also be proved for $\Pi_{J}^{\prime}$ in place of $\Pi^{\prime}$ with those constants independent of $J$.
@@ Acknowledgments
This work was partially supported by the NSFC under grant Nos. 12322116, 12271428, 12326606, 42474139, and the Major projects of the NSFC grants 12090021.

### v2

@@ 6.3. Contraction rates for high-dimensional Besov priors
where $\theta_{lr}=0$ for all $l>J$. Using (1.3) and wavelet characterization of Sobolev norms (see [18, section 4]), we deduce that
@@ 7.1. Some facts about the univariate p-exponential distribution
We first calculate the normalization constant $C_{p,b}$:
@@ 7.2. Concentration of Besov priors
Consider $\Pi^{\prime}$ defined in Definition 2.2. Then,for $p\in[1,2]$, any $h\in B_{pp}^{\alpha}(\mathbb{R}^{d})$ supported on $K$ and any symmetric convex Borel-measurable $A\subset L_{2}(\mathbb{R}^{d})$, it holds
@@
Consider $\Pi^{\prime}_{J}$ defined in Definition 2.3. Then, for $p\geq 1$, any $h\in B_{pp}^{\alpha}(\mathbb{R}^{d})$ supported on $K$ and any symmetric convex Borel-measurable $A\subset L_{2}(\mathbb{R}^{d})$, there exists a constant $c>0$ such that
$\Pi^{\prime}_{J}(P_{J}(h)+A)\geq\big(\frac{1}{2}\big)^{c2^{Jd}}e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{B_{pp}^{\alpha}}}\Pi^{\prime}_{J}(A).$
@@
$\Pi^{\prime}_{J}(A+P_{J}(h))=\int_{A}\exp\Big\{{\sum_{l=-1}^{J}\sum_{r\in R_{l}}\Big(V\big(\frac{\theta_{lr}}{\gamma_{lr}}\big)-V\big(\frac{\theta_{lr}-h_{lr}}{\gamma_{lr}}\big)\Big)}\Big\}\Pi^{\prime}_{J}(d\theta)$
$\geq$ $e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{B_{pp}^{\alpha}}}\int_{A}\exp\Big\{{\sum_{l=-1}^{J}\sum_{r\in R_{l}}\Big(V\big(\frac{\theta_{lr}}{\gamma_{lr}}\big)+V\big(\frac{h_{lr}}{\gamma_{lr}}\big)-V\big(\frac{\theta_{lr}-h_{lr}}{\gamma_{lr}}\big)\Big)}\Big\}\Pi^{\prime}_{J}(d\theta)$
$=$ $e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{B_{pp}^{\alpha}}}\int_{A}\exp\Big\{\sum_{l=-1}^{J-1}\sum_{r\in R_{l}}\Big(V\big(\frac{\theta_{lr}}{\gamma_{lr}}\big)+V\big(\frac{h_{lr}}{\gamma_{lr}}\big)-V\big(\frac{\theta_{lr}-h_{lr}}{\gamma_{lr}}\big)\Big)$
$\qquad\qquad\qquad+\sum_{r\in R_{J}/\{\tilde{r}\}}\Big(V\big(\frac{\theta_{Jr}}{\gamma_{Jr}}\big)+V\big(\frac{h_{Jr}}{\gamma_{Jr}}\big)-V\big(\frac{\theta_{Jr}-h_{Jr}}{\gamma_{Jr}}\big)\Big)\Big\}$
@@
$\Pi^{\prime}_{J}(A+P_{J}(h))$
$\geq$ $e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{B_{pp}^{\alpha}}}\int_{A}\frac{1}{2}\exp\Big\{\sum_{l=-1}^{J-1}\sum_{r\in R_{J}}\Big(V\big(\frac{\theta_{lr}}{\gamma_{lr}}\big)+V\big(\frac{h_{lr}}{\gamma_{lr}}\big)-V\big(\frac{\theta_{lr}-h_{lr}}{\gamma_{lr}}\big)\Big)$
@@
$\geq$ $e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{B_{pp}^{\alpha}}}\int_{A}\big(\frac{1}{2}\big)^{\lvert{R_{J}}\rvert}\exp\Big\{{\sum_{l=-1}^{J-1}\sum_{r\in R_{l}}\Big(V\big(\frac{\theta_{lr}}{\gamma_{lr}}\big)+V\big(\frac{h_{lr}}{\gamma_{lr}}\big)-V\big(\frac{\theta_{lr}-h_{lr}}{\gamma_{lr}}\big)\Big)}\Big\}\Pi^{\prime}_{J}(d\theta)$
$\geq$ $\big(\frac{1}{2}\big)^{\sum_{l=-1}^{J}\lvert{R_{l}}\rvert}e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{B_{pp}^{\alpha}}}\Pi^{\prime}_{J}(A)\geq\big(\frac{1}{2}\big)^{c2^{Jd}}e^{-\frac{1}{p}\lVert{h}\rVert^{p}_{B_{pp}^{\alpha}}}\Pi^{\prime}_{J}(A).$
@@
We note that Lemmas 7.2, 7.5, 7.4, and 7.7 can also be proved for $\Pi_{J}^{\prime}$ in place of $\Pi^{\prime}$ with those constants independent of $J$.
@@ Acknowledgments
This research was partially funded by the National Natural Science Foundation of China (Grant Nos. 12322116, 12271428, 12326606, and 42474139), the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (Grant No. JYB2025XDXM101), the National Key Research and Development Program of China (Grant No. 2022YFA1004100), and the Major Projects of the National Natural Science Foundation of China (Grant Nos. 12090021 and 12090020).